OpenStudy (aworldofcolor):

I was wondering if someone could confirm what I'm doing is right. Situation: You want to buy three books that are on sale at 20% off. The original prices of the books are $2.50, $4.95, and $6.00. How much will you save? We have to do one with distribution and without and I don't know if the distribution one I'm doing is right. I have 2($2.50+$4.95+$6.00) $2.50+$6.00 = $8.50 $8.50+$4.95 = $13.45 0.2*$13.45 = 2.69 and I don't know how to do it without distribution so I was wondering if I could get help on that as well :)

OpenStudy (anonymous):

0.2*13.45=2.69 that is right

OpenStudy (xapproachesinfinity):

correct

OpenStudy (retireed):

I think you got it correct but here is the what I think they are asking 0.2(2.5-4.95+6) = 0.2*2.5+0.2*4.95+0.2*6 = 2.69 2.5 x 0.2 = 0.5 4.95 x 0.2 = 0.99 6 x 0.2 = 1.2 2.69 sum of individual discounts 13.45 x 0.2 = 2.69 sum of books at retail

OpenStudy (aworldofcolor):

Thank all of y'all I got put in the class late and I didn't really learn so I wanted to make sure

OpenStudy (anonymous):

The last answer is correct to find twenty percent of each book remember to multiply by .20